# Gradient-based Sample Selection for Faster Bayesian Optimization

### v1

@@ A.1 Theorem 1:Analysis on subset GP:
We consider a GP scenario where the full GP-UCB at a point $\mathbf{x}$ is given by $\mathrm{UCB}(\mathbf{x})\;=\;\mu(\mathbf{x})\;+\;\beta\,\sigma(\mathbf{x}),$ while its approximate (or sparse) counterpart is $\tilde{\mathrm{UCB}}(\mathbf{x})\;=\;\tilde{\mu}(\mathbf{x})\;+\;\beta\,\tilde{\sigma}(\mathbf{x}).$ We have established the following pointwise error bound:
$\bigl{|}\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr{|}\;\;\leq\;\;\underbrace{\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|\;C_{1}\,(\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}))}_{\displaystyle A}\;+\;\underbrace{\beta\,\|\mathbf{k}_{*\mathcal{D}}\|\sqrt{C_{1}\,(\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2})}}_{\displaystyle\beta B}.$
Hence we obtain the typical form $A=O(\lambda_{M+1}),\quad B=O\bigl{(}\sqrt{\lambda_{M+1}}\bigr{)},$ and $\bigl{|}\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr{|}\;\leq\;A\;+\;\beta\,B.$
We state the theorem that allows merging the term $A+\beta B$ into a single factor $(\beta+\delta)\,\tilde{\sigma}(\mathbf{x})$, under a crucial assumption that $\tilde{\sigma}(\mathbf{x})$ is bounded away from zero.
Theorem A.1 (Single-Penalty Construction).
Let $A\geq 0$ and $B\geq 0$ be constants as above, and assume there exists a lower bound $\tilde{\sigma}(\mathbf{x})\geq\sigma_{\min}>0$ for all $\mathbf{x}$ in the domain. Then one can define $\tilde{\beta}\;=\;\beta\;+\;\delta,\quad\text{where}\quad\delta=\frac{A}{\sigma_{\min}},$ so that $A+\beta\,B\;\;\leq\;\;(\beta+\delta)\,\tilde{\sigma}(\mathbf{x}),$ and therefore $\bigl{|}\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr{|}\;\leq\;(\beta+\delta)\,\tilde{\sigma}(\mathbf{x}).$ Hence, one may write a unified approximate UCB of the form $\widehat{\mathrm{UCB}}(\mathbf{x})\;=\;\tilde{\mu}(\mathbf{x})\;+\;\tilde{\beta}\,\tilde{\sigma}(\mathbf{x}),$ with $\tilde{\beta}:=\beta+\delta$, thereby absorbing both $A$ and $\beta\,B$ into a single penalty term.
Proof.
Step 1: Inequality Setup. We need to ensure that $A+\beta\,B\;\;\leq\;\;(\beta+\delta)\,\tilde{\sigma}(\mathbf{x}),$ for all $\mathbf{x}$ in the domain. By hypothesis, $\tilde{\sigma}(\mathbf{x})\geq\sigma_{\min}$ for all $\mathbf{x}$, hence $(\beta+\delta)\,\tilde{\sigma}(\mathbf{x})\;\;\geq\;\;(\beta+\delta)\,\sigma_{\min}.$ Thus it suffices to impose $A+\beta\,B\;\;\leq\;\;(\beta+\delta)\,\sigma_{\min}.$ Observing that $\beta\,B\leq\beta\,B+A$, it is enough to ensure, individually: $A\;\;\leq\;\;\delta\,\sigma_{\min},\quad\beta\,B\;\;\leq\;\;\beta\,\sigma_{\min}.$ The first part is enforced by $\delta=A/\sigma_{\min}$. The second part can hold if $B\leq\sigma_{\min}$; otherwise, we can slightly adjust $\delta$ to cover that as well.
Step 2: Concluding the Single Penalty. Hence for all $\mathbf{x}$, $A+\beta\,B\;\leq\;\delta\,\sigma_{\min}\;+\;\beta\,\sigma_{\min}\;=\;(\beta+\delta)\,\sigma_{\min}\;\leq\;(\beta+\delta)\,\tilde{\sigma}(\mathbf{x}).$ Therefore, $\bigl{|}\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr{|}\;\leq\;A+\beta\,B\;\leq\;(\beta+\delta)\,\tilde{\sigma}(\mathbf{x}).$ Defining $\tilde{\beta}=\beta+\delta$ yields $\bigl{|}\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr{|}\;\leq\;\tilde{\beta}\,\tilde{\sigma}(\mathbf{x}),$ establishing the desired single-penalty inequality. ∎
Combining Theorem A.1 with our pointwise error bound, we see the following: (1) We have already obtained that $\bigl{|}\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr{|}\;\leq\;A+\beta B,$ where $A=\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|\;C_{1}\,(\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}))$ and $B=\|\mathbf{k}_{*\mathcal{D}}\|\sqrt{C_{1}\,(\lambda_{M+1}+\epsilon\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2})}.$ (2)If, in addition, $\tilde{\sigma}(\mathbf{x})\geq\sigma_{\min}>0$ (i.e. the approximate standard deviation does not vanish), one can pick $\delta\geq A/\sigma_{\min}$ so that $A+\beta\,B\;\leq\;(\beta+\delta)\,\sigma_{\min}\;\leq\;(\beta+\delta)\,\tilde{\sigma}(\mathbf{x}).$ Therefore, the entire error $A+\beta B$ can be encoded by just increasing $\beta$ to $\tilde{\beta}=\beta+\delta$. In practice, however, this tends to be overly conservative in regions where $\tilde{\sigma}(\mathbf{x})$ is large, since $(\beta+\delta)\,\tilde{\sigma}(\mathbf{x})$ might become significantly larger than $A+\beta B$.
In summary, beginning from the Nyström-based UCB error bound
@@
$\bigl{|}\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr{|}\;\leq\;A+\beta B\;\leq\;(\beta+\delta)\,\sigma_{\min}\;\leq\;(\beta+\delta)\,\tilde{\sigma}(\mathbf{x})\quad\bigl{(}A=O(\lambda_{M+1}),\,B=O(\sqrt{\lambda_{M+1}})\bigr{)},$
we have shown that, provided $\tilde{\sigma}(\mathbf{x})\geq\sigma_{\min}>0$, one can inflate $\beta$ to $\beta+\delta$ so as to cover both the linear $A$ term and the $\beta B$ term.
UCB Formulation via a Single Multiplicative Penalty.
Having established that the term $A$ and the $\beta$-proportional term $\beta B$ can be merged under a single inflated parameter $\bigl{|}\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr{|}\;\leq\;A+\beta\,B,$ with constants $A=\|\mathbf{k}_{*\mathcal{D}}\|\|\mathbf{y}\|CC^{\prime}\lambda_{M+1}$ and $B=\beta\|\mathbf{k}_{*\mathcal{D}}\|\sqrt{CC^{\prime}\lambda_{M+1}}$ independent of $\beta$. Assume further that there is a lower bound $\sigma(\mathbf{x})\geq\sigma_{\min}>0$ for all $\mathbf{x}$, so the GP’s true standard deviation never becomes arbitrarily small. In that case, the linear term $A$ and the $\beta$-proportional term $\beta B$ can be combined into a single multiplier by inflating $\beta$ to $\tilde{\beta}\;=\;\beta\;+\;\delta,\quad\text{where}\quad\delta=\frac{A}{\sigma_{\min}}.$ Hence, $A+\beta\,B\;\;\leq\;\;(\beta+\delta)\,\sigma_{\min}\;\;\leq\;\;(\beta+\delta)\,\sigma(\mathbf{x}),$ so that $A+\beta B$ is covered by $(\beta+\delta)\,\sigma(\mathbf{x})$. Thus, in place of the usual $\mu(\mathbf{x})+\beta\,\sigma(\mathbf{x}),$ the new UCB can be written as
$\tilde{\mathrm{UCB}}(\mathbf{x})\;=\;\mu(\mathbf{x})\;+\;\tilde{\beta}\,\sigma(\mathbf{x}),\quad\text{where}\quad\tilde{\beta}=\beta+\delta=\beta+\frac{A}{\sigma_{\min}},\quad B\leq\sigma_{\text{min}}.$
@@ A.2 Theorem 2:Analysis on Regret Bound:
$P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|>L\right)\leq ae^{-L^{2}/b^{2}}\quad\text{for }j=1,\ldots,p.$
@@
$\tilde{\mathrm{UCB}}(\mathbf{x})\;=\;\mu(\mathbf{x})\;+\;(\beta+\frac{A}{\sigma_{\min}})\,\sigma(\mathbf{x}).$
@@
Let $\mathcal{X}\subset[0,r]^{p}$ be compact and convex, $p\in\mathbb{N}$, $r>0$. Under Assumption 1, for any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2p\log\left(n^{2}brp\sqrt{\log\left(\frac{4pa}{\delta}\right)}\right),$i.e.,
$\beta_{n}=2\log\frac{4\pi_{n}}{\delta}+2p\log\left(n^{2}brp\sqrt{\log\left(\frac{4pa}{\delta}\right)}\right)-\frac{A}{\sigma_{\min}},$
where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\quad\pi_{n}>0$. As $n\to\infty$, we obtain a regret bound of $\mathcal{O}^{*}(\sqrt{pN\gamma_{N}})$. Specifically, with $C^{\prime}=\frac{8}{\log(1+\sigma^{-2})},$ we have:
$P\left(R_{N}\leq\sqrt{C^{\prime}N\beta_{N}\gamma_{N}}\right)\geq 1-\delta.$
@@
For any arbitrarily small $\delta_{1}\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{\pi_{n}}{\delta_{1}}$, i.e., $\beta_{n}=2\log\frac{\pi_{n}}{\delta_{1}}-\frac{A}{\sigma_{\min}}$ , where $\sum_{n\geq 1}\pi_{n}^{-1}=1$, $\pi_{n}>0$, then we have
$P\left(|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})|\leq d\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\delta$
@@
Assuming we are at stage $n$, all past decisions $\mathbf{x}_{1:n-1}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n-1}\}$ made after the initial design are deterministic given $\mathbf{y}_{1:n-1}=\{y_{1},\ldots,y_{n-1}\}$. For any $\mathbf{x}_{n}\in\mathbb{R}^{p}$, we have $f(\mathbf{x}_{n})\sim\mathcal{N}(\mu_{n-1}(\mathbf{x}_{n}),\sigma^{2}_{n-1}(\mathbf{x}_{n}))$.
For a standard normal variable $r\sim\mathcal{N}(0,1)$, the probability of being above a certain constant $c$ is written as:
@@
[…]
$=e^{-c^{2}/2}P(r>0)$
$=\frac{1}{2}e^{-c^{2}/2}$
where the inequality holds due to the fact that $e^{-c(r-c)}\leq 1$ for $r\geq c>0$.
Plugging in $r=\frac{f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})}{\sigma_{n-1}(\mathbf{x}_{n})}$ and $c=\tilde{\beta_{n}}^{1/2}$, we have:
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|>\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\leq e^{-\frac{\tilde{\beta_{n}}}{2}}.$
@@
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-e^{-\frac{\tilde{\beta_{n}}}{2}}.$
@@
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\sum_{n\geq 1}\frac{\delta}{\pi_{n}}=1-\delta.$
@@
For any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{|\mathcal{X}_{n}|\pi_{n}}{\delta},i.e.,\beta_{n}=2\log\frac{|\mathcal{X}_{n}|\pi_{n}}{\delta}-\frac{A}{\sigma_{\min}},$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\quad\pi_{n}>0,$ then we have
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\delta\quad\text{for }\forall\mathbf{x}_{n}\in\mathcal{X}_{n},\forall n\geq 1.$
@@
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-e^{-\frac{\tilde{\beta_{n}}}{2}}.$
@@
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-|\mathcal{X}_{n}|e^{-\frac{\tilde{\beta_{n}}}{2}},\quad\forall\mathbf{x}_{n}\in\mathcal{X}_{n}.$
@@
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\sum_{n\geq 1}\frac{\delta}{\pi_{n}}=1-\delta,$
@@
For any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{2\pi_{n}}{\delta}+2p\log\left(n^{2}brp\sqrt{\log\left(\frac{2pa}{\delta}\right)}\right),i.e.,\beta_{n}=2\log\frac{2\pi_{n}}{\delta}+2p\log\left(n^{2}brp\sqrt{\log\left(\frac{2pa}{\delta}\right)}\right)-\frac{A}{\sigma_{\min}},$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1$, $\pi_{n}>0$, $p\in\mathbb{N}$ is the dimensionality of the feature space, and $r>0$ is the length of the domain in a compact and convex set $\mathcal{X}\subset[0,r]^{p}$. Given constants $a$, $b$ and $L$, assume that the kernel function $k(\mathbf{x},\mathbf{x}^{\prime})$ satisfies the following Lipschitz continuity for the confidence bound of the derivatives of GP sample paths $f$:
$P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|>L\right)\leq ae^{-L^{2}/b^{2}},\quad j=1,\dots,p,$
@@
$P\left(\left|f(\mathbf{x}^{*})-\mu_{n-1}([\mathbf{x}^{*}]_{n})\right|\leq d\tilde{\beta_{n}^{\frac{1}{2}}}\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}\right)\geq 1-\delta,\quad\forall n\geq 1.$
@@
$P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|<L\right)\geq 1-pae^{-L^{2}/b^{2}}$
@@
$P\left(\left|f(\mathbf{x})-f(\mathbf{x}^{\prime})\right|\leq L\|\mathbf{x}-\mathbf{x}^{\prime}\|_{1}\right)\geq 1-pae^{-L^{2}/b^{2}}\quad\forall\mathbf{x}\in\mathcal{X}$
which is a confidence bound that applies to $\mathbf{x}^{*}$ as well. For a discretization $\mathcal{X}_{n}$ of size $(\tau_{n})^{p}$, i.e., each coordinate space of $\mathcal{X}_{n}$ has a total of $\tau_{n}$ discrete points, we have the following bound on the closest point $[\mathbf{x}]_{n}$ to $\mathbf{x}$ in $\mathcal{X}_{n}$ to ensure a dense set of discretizations:
$\|\mathbf{x}-[\mathbf{x}]_{n}\|_{1}\leq\frac{rp}{\tau_{n}}.$
Now, setting $pae^{-L^{2}/b^{2}}=\frac{\delta}{2}$, i.e., $L=b\sqrt{\log\left(\frac{2pa}{\delta}\right)}$, gives the following:
$P\left(\left|f(\mathbf{x})-f(\mathbf{x}^{\prime})\right|\leq b\sqrt{\log\left(\frac{2pa}{\delta}\right)}\|\mathbf{x}-\mathbf{x}^{\prime}\|_{1}\right)\geq 1-\frac{\delta}{2}\quad\forall\mathbf{x}\in\mathcal{X}.$

### v2

@@ A.1 Theorem 1:Analysis on subset GP:
We consider a GP scenario where the full GP-UCB at a point $\mathbf{x}$ is given by $\mathrm{UCB}(\mathbf{x})=\;\mu(\mathbf{x})+\;\beta_{n}\,\sigma(\mathbf{x}),$ while its approximate counterpart is $\tilde{\mathrm{UCB}}(\mathbf{x})=\;\tilde{\mu}(\mathbf{x})+\;\beta_{n}\,\tilde{\sigma}(\mathbf{x}).$ We have established the following pointwise error bound: $\bigl{|}\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr{|}\;\leq\;A_{n}+\beta_{n}\,B_{n},$ where
$A_{n}=\|\mathbf{k}_{*\mathcal{D}}\|\;\|\mathbf{y}\|\;C_{M}\;\Bigl{(}\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\Bigr{)},~{}B_{n}=\|\mathbf{k}_{*\mathcal{D}}\|\;\sqrt{\,C_{M}\;\Bigl{(}\lambda_{M+1}+\epsilon_{g}\sum_{i=1}^{N}(\mathbf{K}_{\mathcal{D}\mathcal{D}})_{ii}^{2}\Bigr{)}}.$ (17)
Here, $A_{n},B_{n}$ are determined by the number of iterations $n$, the same below. We now absorb both terms into a single, $n$–dependent penalty. Define $\sigma_{\min}(n)\;=\;\min_{x\in\mathcal{X}}\,{\sigma}_{n}(x),\qquad\delta_{n}\;=\;\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)},\qquad\tilde{\beta}_{n}\;=\;\beta_{n}+\delta_{n}.$ Since for all $x$, ${\sigma}_{n}(x)\geq\sigma_{\min}(n)>0$, it follows that $A_{n}+\beta_{n}\,B_{n}=\;\delta_{n}\,\sigma_{\min}(n)\;\leq\;\delta_{n}\,{\sigma}_{n}(x),$ and hence
@@
$\bigl{|}\mathrm{UCB}(\mathbf{x})-\tilde{\mathrm{UCB}}(\mathbf{x})\bigr{|}\;\leq\;A_{n}+\beta_{n}\,B_{n}\;\leq\;\delta_{n}\,{\sigma}_{n}(\mathbf{x})=\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)}{\sigma}_{n}(\mathbf{x}).$ (18)
Therefore, we may write the unified approximate UCB as
$\tilde{\mathrm{UCB}}(\mathbf{x})={\mu}(\mathbf{x})+\beta_{n}\,{\sigma}_{n}(\mathbf{x})+\delta_{n}\,{\sigma}_{n}(x)=\;{\mu}(\mathbf{x})+\;\tilde{\beta}_{n}\,{\sigma}_{n}(\mathbf{x}),$ (19)
where $\tilde{\beta}_{n}=\;\beta_{n}+\;\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)}.$ This completes the single-penalty construction.
@@ A.2 Theorem 2:Analysis on Regret Bound:
$P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|>L\right)\leq ae^{-L^{2}/b^{2}}\quad\text{for }j=1,\ldots,d.$ (20)
@@
$\tilde{\mathrm{UCB}}(\mathbf{x})\;=\;\mu(\mathbf{x})\;+\;(\beta_{n}+\frac{A_{n}+\beta_{n}\,B_{n}}{\sigma_{\min}(n)})\,\sigma(\mathbf{x}).$ (21)
@@
Let $\mathcal{X}\subset[0,r]^{d}$ be compact and convex, $d\in\mathbb{N}$, $r>0$. Under Assumption 1, for any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{4\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{4da}{\delta}\right)}\right),$i.e.,
[…]
where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\quad\pi_{n}>0$. As $n\to\infty$, we obtain a regret bound of $\mathcal{O}^{*}(\sqrt{d{T_{N}}\gamma_{T_{N}}})$. Specifically, with $C_{1}=\frac{8}{\log(1+\sigma^{-2})},$ we have:
$P\left(R_{T_{N}}\leq\sqrt{C_{1}{T_{N}}\beta_{{T_{N}}}\gamma_{T_{N}}}\right)\geq 1-\delta.$ (23)
@@
For any arbitrarily small $\delta_{1}\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{\pi_{n}}{\delta_{1}}$, i.e., $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,(2\log\frac{\pi_{n}}{\delta_{1}})\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ , where $\sum_{n\geq 1}\pi_{n}^{-1}=1$, $\pi_{n}>0$, then we have
$P\left(|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})|\leq d\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\delta$ (24)
@@
Assuming we are at stage $n$, all past decisions $\mathbf{x}_{1:n-1}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n-1}\}$ made after the initial design are deterministic given $\mathbf{y}_{1:n-1}=\{y_{1},\ldots,y_{n-1}\}$. For any $\mathbf{x}_{n}\in\mathbb{R}^{d}$, we have $f(\mathbf{x}_{n})\sim\mathcal{N}(\mu_{n-1}(\mathbf{x}_{n}),\sigma^{2}_{n-1}(\mathbf{x}_{n}))$. For a standard normal variable $r\sim\mathcal{N}(0,1)$, the probability of being above a certain constant $c$ is written as:
@@
where the inequality holds due to the fact that $e^{-c(r-c)}\leq 1$ for $r\geq c>0$. Plugging in $r=\frac{f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})}{\sigma_{n-1}(\mathbf{x}_{n})}$ and $c=\tilde{\beta_{n}}^{1/2}$, we have:
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|>\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\leq e^{-\frac{\tilde{\beta_{n}}}{2}}.$ (26)
@@
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-e^{-\frac{\tilde{\beta_{n}}}{2}}.$ (27)
@@
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{1/2}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\sum_{n\geq 1}\frac{\delta}{\pi_{n}}=1-\delta.$ (28)
@@
For any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{|\mathcal{X}_{n}|\pi_{n}}{\delta},i.e.,$ $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,(2\log\frac{|\mathcal{X}_{n}|\pi_{n}}{\delta})\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1,\quad\pi_{n}>0,$ then we have
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\delta\quad\text{for }\forall\mathbf{x}_{n}\in\mathcal{X}_{n},\forall n\geq 1.$ (29)
@@
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-e^{-\frac{\tilde{\beta_{n}}}{2}}.$ (30)
@@
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-|\mathcal{X}_{n}|e^{-\frac{\tilde{\beta_{n}}}{2}},\quad\forall\mathbf{x}_{n}\in\mathcal{X}_{n}.$ (31)
@@
$P\left(\left|f(\mathbf{x}_{n})-\mu_{n-1}(\mathbf{x}_{n})\right|\leq\tilde{\beta_{n}}^{\frac{1}{2}}\sigma_{n-1}(\mathbf{x}_{n})\right)\geq 1-\sum_{n\geq 1}\frac{\delta}{\pi_{n}}=1-\delta,$ (32)
@@
For any arbitrarily small $\delta\in(0,1)$, choose $\tilde{\beta_{n}}=2\log\frac{2\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{2da}{\delta}\right)}\right),i.e.,$ $\beta_{n}\;=\;\frac{\sigma_{\min}(n)\,\Bigl{[}2\log\frac{2\pi_{n}}{\delta}+2d\log\left(n^{2}brd\sqrt{\log\left(\frac{2da}{\delta}\right)}\right)\Bigr{]}\;-\;A_{n}}{\sigma_{\min}(n)\;+\;B_{n}}\,.$ where $\sum_{n\geq 1}\pi_{n}^{-1}=1$, $\pi_{n}>0$, $d\in\mathbb{N}$ is the dimensionality of the feature space, and $r>0$ is the length of the domain in a compact and convex set $\mathcal{X}\subset[0,r]^{d}$. Given constants $a$, $b$ and $L$, assume that the kernel function $k(\mathbf{x},\mathbf{x}^{\prime})$ satisfies the following Lipschitz continuity for the confidence bound of the derivatives of GP sample paths $f$:
$P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|>L\right)\leq ae^{-L^{2}/b^{2}},\quad j=1,\dots,d,$ (33)
@@
$P\left(\left|f(\mathbf{x}^{*})-\mu_{n-1}([\mathbf{x}^{*}]_{n})\right|\leq d\tilde{\beta_{n}^{\frac{1}{2}}}\sigma_{n-1}([\mathbf{x}^{*}]_{n})+\frac{1}{n^{2}}\right)\geq 1-\delta,\quad\forall n\geq 1.$ (34)
@@
$P\left(\sup_{\mathbf{x}\in\mathcal{X}}\left|\frac{\partial f}{\partial x_{j}}\right|<L\right)\geq 1-dae^{-L^{2}/b^{2}}$ (35)
@@
$P\left(\left|f(\mathbf{x})-f(\mathbf{x}^{\prime})\right|\leq L\|\mathbf{x}-\mathbf{x}^{\prime}\|_{1}\right)\geq 1-dae^{-L^{2}/b^{2}}\quad\forall\mathbf{x}\in\mathcal{X}$ (36)
which is a confidence bound that applies to $\mathbf{x}^{*}$ as well. For a discretization $\mathcal{X}_{n}$ of size $(\tau_{n})^{d}$, i.e., each coordinate space of $\mathcal{X}_{n}$ has a total of $\tau_{n}$ discrete points, we have the following bound on the closest point $[\mathbf{x}]_{n}$ to $\mathbf{x}$ in $\mathcal{X}_{n}$ to ensure a dense set of discretizations:
$\|\mathbf{x}-[\mathbf{x}]_{n}\|_{1}\leq\frac{rd}{\tau_{n}}.$ (37)
Now, setting $dae^{-L^{2}/b^{2}}=\frac{\delta}{2}$, i.e., $L=b\sqrt{\log\left(\frac{2da}{\delta}\right)}$, gives the following:
$P\left(\left|f(\mathbf{x})-f(\mathbf{x}^{\prime})\right|\leq b\sqrt{\log\left(\frac{2da}{\delta}\right)}\|\mathbf{x}-\mathbf{x}^{\prime}\|_{1}\right)\geq 1-\frac{\delta}{2}\quad\forall\mathbf{x}\in\mathcal{X}.$ (38)